One particular part is used approximately 2,000,000 per year. It is produced in batches of 2,000,000. Recently, you have been put in charge of determining a new strategy and look at batch sizes of 50,000. You determine that material cost is $1.25, labor cost $1.50 and distribution cost $0.25 per part. An interest rate of 22% has been established to represent opportunity and storage costs. a) What is the cost of the first strategy? b) What is the optimal EOQ? c) Suppose your plant can make a maximum of 75,000 parts a week. What is the optimal size production run? Ordering cost is $100.
Answer #1

cost of manufacturing = 2000000x(1.25+1.5+0.25) = 6000000

cost of holding the inventory = avg inventory x cost of holding one unit per year

= 2000000/2 x 0.25 = 250000

cost of ordering = 1 order x 100 = 100

total cost of first strategy = 6250100

Optimal EOQ = [ 2 x annual demand x cost of ordering / cost of holding one unit for a year]1/2

= [ 2 x 2000000x100 /0.25]1/2 = 40000

demand per week = 2000000 /52 = 38461

Demand / production ratio = 38461/75000 = 0.5128

Optimum production quantity = [ 2 x annual demand x order cost / cost of holding one unit x ( 1-d/p)]1/2

= [ 2 x 2000000 x 100 / 0.25 x( 1-0.5128)]1/2

= 57309